A wire with a resistance of 5Ω is drawn out through a die so that its new length is four times its original length. What is the resistance of the longer wire when the resistivity and density of the material are unchanged?


1
Expert's answer
2022-04-11T12:44:22-0400
"R=\rho \frac{l}{A}"

The density of the wire remains the same, hence the volume of the wire would be unchanged

"V=lA=const"

The new resistance of the wire

"R'=\rho'\frac{l'}{A'}=\rho\frac{4l}{A/4}=16R\\
=16*5\:\Omega=80\:\Omega"